OpenStudy (anonymous):

Solve the solution set using interval notation. 1.1x- 3- 0.6x _>_ 15

OpenStudy (anonymous):

\[\ge\] It is that

OpenStudy (anonymous):

(11/10)x-3-(6/10)x >= 15 (5/10)x -3 >=15 (1/2)x -3 >= 15 x-6>=30 x>=36 [36,infinity)

OpenStudy (anonymous):

How do you get 10?

OpenStudy (anonymous):

1.1=(11/10) 0.6=(6/10)